# Bayesian Optimization for Non-Convex Two-Stage Stochastic Optimization Problems

### v1

@@ Appendix C Optimization of the acquisition functions using multi-start L-BFGS-B
The acquisition functions for joint, alternating and two-step KG in Equations 12, 17, 15c and 19 are optimized using multi-start L-BFGS using the heuristics implemented in the function optimize_acqf in BoTorch (Balandat et al., 2020). The function is called with nonnegative=True since knowledge gradient is non-negative (Proposition A.4).
@@
The start points for L-BFGS are generated by first evaluating the acquisition function at 256 locations in $\tilde{{\mathbb{X}}}_{\text{raw}}\subset\tilde{{\mathbb{X}}}$ determined by a scrambled Sobol’ sequence. These are referred to as ‘raw samples’. The raw samples are subsampled to those at least 0.01% of the maximum, $\alpha_{\text{max}}=\max_{\tilde{{\bm{x}}}\in\tilde{{\mathbb{X}}}_{\text{raw}}}\alpha(\tilde{{\bm{x}}})$, to remove values close to zero. This gives, $\tilde{{\mathbb{X}}}_{\text{raw},\gg 0}=\{\tilde{{\bm{x}}}\in\tilde{{\mathbb{X}}}_{\text{raw}}\,:\,\alpha(\tilde{{\bm{x}}})>10^{-4}\alpha_{\text{max}}\}$. Finally, 10 restart locations are chosen from $\tilde{{\mathbb{X}}}_{\text{raw},\gg 0}$ using Boltzmann sampling. That is, a subsample is taken independently, without replacement using probabilities
@@ Appendix D Optimization for the final recommendation
Knowledge gradient acquisition functions are one-step Bayes-optimal when the final recommendation is given by the maximum of the posterior mean. For the joint and alternating knowledge gradient, this is given by Equation 10. The optimization is done in two stages. First, the optimal fixed design, $\hat{{\bm{x}}}^{*n}$, is chosen to maximize Equation 13 using multi-start L-BFGS. Then, the optimal control policy, $\hat{{\bm{g}}}^{*n}$, is optimized for each ${\bm{u}}$ of interest using single-start L-BFGS. This section explains in detail how these optimizations are performed.
@@ D.1 One-shot optimization of fixed design via multi-start L-BFGS-B
The optimal fixed design, $\hat{{\bm{x}}}^{*n}\in{\mathbb{X}}$, is found by rewriting Equation 13 as
@@
Batch limit for raw samples $>N_{\text{raw},y}$
@@ D.2 Single-start L-BFGS-B for optimization of control policy
Once $\hat{{\bm{x}}}^{*n}$ has been found, we must optimize to find $\hat{{\bm{g}}}^{*n}({\bm{u}})\in{\mathbb{Y}}$ for the ${\bm{u}}\in{\mathbb{U}}$ of interest. For example, when computing the simple regret from Equation 21, we wish to compute $\hat{{\bm{g}}}^{*n}({\bm{u}}^{j})$ for a sample ${\bm{u}}^{1},\dots,{\bm{u}}^{N_{u}}$ of the environmental variable ${\mathbf{u}}$, independent of the one used to optimize $\hat{{\bm{x}}}^{*n}$. Empirically, we found that a single start L-BFGS-B is sufficient for this problem. Again, we use optimize_acqf from BoTorch, with $N_{\text{raw},y}=32$ or $128$ raw samples depending on the problem dimension, as in Section D.1. The best of the raw samples is used as the initial point for the L-BFGS-B. Optimizations for the different ${\bm{u}}$ are run separately.
@@ E.1 Experimental parameters
Table 4 summarizes the parameters used for the optimization of the acquisition function and estimation of the the (average) simple regret in the empirical studies. These are in addition to the parameters which have already been specified in Tables 1, 2 and 3.
@@
E.2 Optical table test problem
@@
E.3 Fitting the GP Surrogate Model
@@
The hyperparameters are updated at the start of each iteration, to incorporate the new observations. The first time the GP is fitted, the initial observations are normalized to have zero mean and unit variance and MAP estimates are generated for all hyperparameters to fit the normalized observations. On future iterations, the data is renormalized and the constant mean from the initial fit is updated to the new normalization. New MAP estimates for all hyperparameters except the constant mean are found. When using the GP in the acquisition function or when making recommendations, the GP is untransformed. That is, if $f_{\text{norm}}\sim\mathcal{GP}(\mu_{\text{norm}},k_{\text{norm}})$ is the GP fitted to the normalized data, then $f=a+bf_{\text{norm}}$ is the GP referred to in the main text, where $a$ and $b$ are the mean and sample standard deviation of the observations.

### v2

@@ Appendix C Optimization of the acquisition functions using multi-start L-BFGS-B
The acquisition functions for joint, alternating and two-step KG in Equations 12, 17, 15c and 19 are optimized using multi-start L-BFGS-B using the heuristics implemented in the function optimize_acqf in BoTorch (Balandat et al., 2020). The function is called with nonnegative=True since knowledge gradient is non-negative (Proposition A.4).
@@
The start points for L-BFGS-B are generated by first evaluating the acquisition function at 256 locations in $\tilde{{\mathbb{X}}}_{\text{raw}}\subset\tilde{{\mathbb{X}}}$ determined by a scrambled Sobol’ sequence. These are referred to as ‘raw samples’. The raw samples are subsampled to those at least 0.01% of the maximum, $\alpha_{\text{max}}=\max_{\tilde{{\bm{x}}}\in\tilde{{\mathbb{X}}}_{\text{raw}}}\alpha(\tilde{{\bm{x}}})$, to remove values close to zero. This gives, $\tilde{{\mathbb{X}}}_{\text{raw},\gg 0}=\{\tilde{{\bm{x}}}\in\tilde{{\mathbb{X}}}_{\text{raw}}\,:\,\alpha(\tilde{{\bm{x}}})\geq 10^{-4}\alpha_{\text{max}}\}$. Finally, 10 restart locations are chosen from $\tilde{{\mathbb{X}}}_{\text{raw},\gg 0}$ using Boltzmann sampling. That is, a subsample is taken independently, without replacement using probabilities
@@
There are some differences to this approach for the supply chain problem, which are detailed separately in Section E.3.2.
@@ Appendix D Optimization for the final recommendation
Knowledge gradient acquisition functions are one-step Bayes-optimal when the final recommendation is given by the maximum of the posterior mean. For the joint and alternating knowledge gradient, this is given by Equation 10. The optimization is done in two stages. First, the optimal fixed design, $\hat{{\bm{x}}}^{*n}$, is chosen to maximize Equation 13 using multi-start L-BFGS-B. Then, the optimal control policy, $\hat{{\bm{g}}}^{*n}$, is optimized for each ${\bm{u}}$ of interest using single-start L-BFGS-B. This section explains in detail how these optimizations are performed.
Note that this approach is not used for the supply chain problem, where instead recommendations are generated using an exhaustive search following (Xie et al., 2021).
@@ D.1 One-shot optimization of fixed design via multi-start L-BFGS-B
The optimal fixed design is approximated by rewriting Equation 13 as
@@
Batch limit for raw samples $2^{17}$
@@ D.2 Single-start L-BFGS-B for optimization of control policy
Once the recommended fixed design $\hat{{\bm{x}}}^{*n}$ has been found, we must optimize to find the recommended control policy $\hat{{\bm{g}}}^{*n}({\bm{u}})\in{\mathbb{Y}}$ for the ${\bm{u}}\in{\mathbb{U}}$ of interest. For example, when computing the simple regret from Equation 21, we wish to compute $\hat{{\bm{g}}}^{*n}({\bm{u}}^{j})$ for a sample ${\bm{u}}^{1},\dots,{\bm{u}}^{N_{u}}$ of the environmental variable ${\mathbf{u}}$, independent of the one used to optimize $\hat{{\bm{x}}}^{*n}$. Empirically, we found that a single start L-BFGS-B is sufficient for this problem. Again, we use optimize_acqf from BoTorch, with $N_{\text{raw},y}=32$ or $128$ raw samples depending on the problem dimension, as in Section D.1. The best of the raw samples is used as the initial point for the L-BFGS-B. Optimizations for the different ${\bm{u}}$ are run separately (corresponding to a batch_limit of 1).
@@ E.1 Experimental parameters
Table 5 summarizes the parameters used for the optimization of the acquisition function and estimation of the the (average) simple regret in the empirical studies. These are in addition to the parameters which have already been specified in Tables 2, 3 and 4.
@@
For all GP generated test problems, the fixed design used by the two-step algorithm in the first step is ${\bm{x}}_{\text{step-1}}=(0.5,\dots,0.5)\in\mathbb{R}^{d_{x}}$.
For the optical table problem, the fixed design is the spring constant $k$, the adjustable variable is the damping coefficient $c$, and the uncertain environmental parameter is the angular frequency of the floor vibration $\omega$. The search space for the optimization is $12\,N\,mm^{-1}\leq k\leq 50\,N\,mm^{-1}$, $1\,Ns\,mm^{-1}\leq c\leq 10\,Ns\,mm^{-1}$ and $1\,Hz\leq\omega/2\pi\leq 100\,Hz$. The fixed design used in the first step of the two-step algorithms is $k_{\text{step-1}}=31\,N\,mm^{-1}$.
@@
E.2 Analytical solution for the optical table test problem
@@
E.3 Supply chain management problem
The supply chain problem has several features not considered in the other test problems. Namely, it has normally distributed environmental variables, it has constraints coupling the fixed and adjustable variables, and as defined in (Xie et al., 2021) it has a finite search space (${\bm{x}}$ and ${\bm{y}}$ can only take integer values and ${\bm{x}}$ must be a multiple of $20$). These lead to a some natural modifications to the implementation of the algorithm.
E.3.1 Normally distributed environmental variables
The approximations to the knowledge gradient acquisition functions in Equations 12, 15c, 17 and 19 are defined using an MC or qMC estimate of the expectation $\mathbb{E}_{\mathbf{u}}[\cdot]$ over the environmental variables. Since qMC estimates are lower variance unbiased estimators, we continue to use a qMC estimate for the expectation over the normally distributed environmental variables ${\mathbf{u}}$. Specifically, we generate the sample by first generating a Sobol’ sample of $N_{u}$ points in $[0,1]^{d_{u}}$, and then pass these point-wise and coordinate-wise through the inverse CDF $F_{u}^{-1}$ of the normal distribution. That is,
$({\bm{z}}^{1},\dots,{\bm{z}}^{N_{u}})\sim\mathrm{Sobol}([0,1]^{d_{u}},N_{u}),$
$F_{u}(u^{\prime})=F_{u,j}(u^{\prime})=\mathbb{P}({\textnormal{u}}_{j}\leq u^{\prime})\quad\forall\,j,$
$\forall\,i=1,\dots,N_{u}\,\forall\,j=1,\dots,4\quad{u}_{j}^{i}=F_{u}^{-1}({z}_{j}^{i}).$
We use the implementation in BoTorch (Balandat et al., 2020) and use the same method for the qMC estimates over ${\mathbf{u}}$ when generating recommendations and estimating regret.
In addition to requiring a qMC estimate of the expectation over the normal environmental variables, we also choose to use normally distributed ${\mathbf{u}}$ in the initial sample. For the other test problems the initial sample is a Sobol’ sample in $[0,1]^{d_{x}+d_{y}+d_{u}}$, rescaled to the bounds of the test problem in the case of the optical table. For the supply chain problem, we instead independently generate a uniform Sobol’ sample on $[0,1]^{d_{x}+d_{y}}$ for ${\mathbb{X}}\times{\mathbb{Y}}$ and a normal Sobol’ sample on $\mathbb{R}^{d}_{u}$ for ${\mathbb{U}}$. These are concatenated to give a sample for $[0,1]^{d_{x}+d_{y}}\times{\mathbb{R}}^{d_{u}}$. Since the problem has constraints, we discard infeasible points from the initial design and continue to generate samples using the Sobol’ sequences until sufficiently many feasible initial points have been sampled.
Finally, we observe that, unlike uniformly distributed environmental variables, normally distributed ${\mathbf{u}}$ have no natural bounds. When optimizing the acquisition function, we use the $1\%$ and $99\%$ marginal quantiles to give bounds on ${\mathbb{U}}$. When normalizing the input space before modeling with a GP, we use a box with edges one standard deviation either side of the mean. This is preferable to using the $1\%$ and $99\%$ quantiles which were used for the bounds, because the latter would concentrate the majority of probability mass of ${\mathbf{u}}$ in a very small volume relative to the priors used for the length scales of the GP.
E.3.2 Optimization of the acquisition function
The usual process we have followed for optimizing the acquisition function is explained in Appendix C. We first generate promising initial samples by evaluating the acquisition function at a large number of points, then selecting a random sample of the best with Boltzmann sampling, always ensuring the best evaluated point is among the chosen ones. Finally, we optimize each of these using L-BFGS-B and select the best.
However, for the supply chain problem we must satisfy the constraint $y_{1}\leq x/20$, we must restrict to only 10 different $(s,S)$-ordering policies for the raw chemical, and we must ensure the soy $x$ and daily production $y_{1}$ are chosen from among the finite set detailed in Table 1 of the main text. We achieve this with the following modifications to the process in Appendix C. First, when generating ‘raw samples’, we use the MCMC approach implemented in BoTorch (Balandat et al., 2020) to generate a sample for over all but the $(s,S)$ variables. We then take the Cartesian product with the 10 possible values of $(s,S)$ and proceed with the usual Boltzmann sampling to choose promising candidates for the 10 restarts, again ensuring the best of the raw samples is one of the restarts. Instead of L-BFGS-B, we use the SLSQP implementation in scipy to optimize these restarts subject to the constraint $y_{1}\leq x/20$. We keep $(s,S)$ fixed during this optimization. Finally, the best of these optimized candidates is rounded to give the next sample location as follows. First the soy quantity $x$ is rounded to the nearest 20, then the daily production $y_{1}$ is clipped to the nearest integer in the range $[0,x/20]$.
E.3.3 Modifications to the two-step algorithms
The two-step knowledge gradient and random sampling algorithms generate a control policy in the first step which is then held fixed in the second step. For our other experiments, this policy is evaluated by running an optimization using L-BFGS-B to find the maximum of the final posterior mean from step 1. For the supply chain problem, we instead use an exhaustive search to choose the best ${\mathbf{y}}=(y_{1},s,S)$ for each ${\bm{u}}$.
However, the soy $x$ is optimized in the second step, which affects the feasible region for $y_{1}$. For convenience of the implementation, if the chosen $y_{1}$ lies outside the feasible region, then we simply round it to the nearest integer in $[0,x/20]$.
E.4 Fitting the GP Surrogate Model
@@
The hyperparameters are updated at the start of each iteration, to incorporate the new observations. The data is normalized to have zero mean and unit variance, and MAP estimates are generated for all hyperparameters to fit the normalized observations. When using the GP in the acquisition function or when making recommendations, the GP is untransformed. That is, if $f_{\text{norm}}\sim\mathcal{GP}(\mu_{\text{norm}},k_{\text{norm}})$ is the GP fitted to the normalized data, then $f=a+bf_{\text{norm}}$ is the GP referred to in the main text, where $a$ and $b$ are the mean and sample standard deviation of the observations.